Differentiate between the terms revenue and profit. Assume that a firm sells 20 units of a good at a price of $4 per unit. If the average total cost of the firm is $2.50 per unit, calculate the firm’s profit.
1
Expert's answer
2020-05-17T17:47:25-0400

Revenue= Price * Quantity

20 * 4= $80

Profit = Revenue - Cost

Cost = Average Cost * Quantity

2.5 * 20 = $50

Profit = 80 - 50 = $30

Profit = $30